OpenStudy (anonymous):

−48 ÷ (7 − 9)^8-2[1 − 5(2 − 6) + 3^2] simplify and show work is the answer 968/12

OpenStudy (anonymous):

-48/(-2)^8 -2[1+20+9] =-48/256-60 =-3/16-60 -963/16